# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ B Predator-Prey Model
We consider a predator-prey system with a minimalistic attraction-repulsion mechanism, such as dogs-sheeps. Thus, the microscopic interaction rules are $v_{ds}=-1$, $v_{sd}=1$, $v_{dd}=0$, $v_{ss}=-1$ which implies $v_{ds}^{r}=-1$, $v_{sd}^{r}=1$, $v_{dd}^{r}=0$, $v_{ss}^{r}=-1$. This leads to the macroscopic interaction rules, $V_{sd}^{r}=\frac{\beta^{2}}{2\Omega}$, $V_{sd}^{nr}=\beta,V_{ds}^{nr}=-\beta$, and $V_{ss}^{r}=-\beta+\frac{\beta^{2}}{2\Omega}$. We have intentionally chosen $v_{ss}=-1$, which models an attractive interaction between sheep, physically corresponding to the flocking of the sheep herd. The effective coarse-grained energy functional for this model reads:
@@
where, the reciprocal macroscopic Boltzmann weights are $\mu_{d}^{r}=\ln{\rho_{d}}+V_{sd}^{r}\rho_{s}$ and $\mu_{s}^{r}=\ln{\rho_{s}}+V_{sd}^{r}\rho_{d}+V_{ss}^{r}\rho_{s}$, and the non-reciprocal macroscopic Boltzmann weights are $F_{d}^{nr}=V_{ds}^{nr}\rho_{s}$, $F_{s}^{nr}=V_{sd}^{nr}\rho_{d}$. The diffusive mobilities are given by $D_{d}=\tilde{d}_{d}\rho_{d}e^{(V_{ds}^{r}+V_{ds}^{nr})\rho_{s}}$ and $D_{s}=\tilde{d}_{s}\rho_{s}e^{\left(V_{ss}^{r}\rho_{s}+(V_{sd}^{r}+V_{sd}^{nr})\rho_{d}\right)}$. $F_{ds}^{ch}=\vec{F}_{i}^{sp}=0$, thus $\langle\dot{\Sigma}^{sp}\rangle=\langle\dot{\Sigma}_{\mathcal{R}}^{nr}\rangle=\langle\dot{\Sigma}^{ch}\rangle=0$. The expression for $\langle\dot{\Sigma}^{nr}\rangle$ reads:
@@
The second law of thermodynamics is $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+d_{t}S^{gb}+\langle\dot{\Sigma}^{nr}\rangle$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle$.
The change in the symmetric part of macrostate correlations is $O(1/\Omega)$, as $V_{sd}^{nr}\propto O(1/\Omega)$, compared to section A, where $V_{sd}^{nr}\propto O(1)$, which highlights an importance difference between two models. Importantly, despite the similarity of $\langle\dot{\Sigma}^{nr}\rangle$ for both models, the symmetric part of macrostate correlations distinguishes between the underlying different physical mechanisms that generate the same vorticity currents. In section A, it is one-way attraction, and here it is a mutual attraction-repulsion mechanism that leads to non-reciprocal vorticity currents.
@@ A Single species
The Active Ising model consists of two different types of self-propelled particles, positive and negative [68]. The particles of the same type attract each otherwise repeals. Hence, microscopic interaction rules are $v_{++}=-1$, $v_{--}=-1$ and $v_{-+}=1$. This leads to the macroscopic interaction coefficients $V_{--}^{r}=V_{++}^{r}=-1+\frac{\beta}{2\Omega}$ and $V_{-+}^{r}=1+\frac{\beta}{2\Omega}$. Here, we consider the thermodynamically consistent Active Ising model. The macroscopic reciprocal Boltzmann weights are $\mu_{+}^{r}=\log{\rho_{+}}+\beta\left(V_{++}^{r}\rho_{+}+V_{+-}^{r}\rho_{-}\right)$ and $\mu_{-}^{r}=\log{\rho_{-}}+\beta\left(V_{--}^{r}\rho_{-}+V_{-+}^{r}\rho_{+}\right)$. The diffusive mobility reads $D_{+}=\tilde{d}e^{\mu_{+}^{r}},D_{-}=\tilde{d}e^{\mu_{-}^{r}}$. We report a detailed study in Ref. [72, 86] and outline the thermodynamic consequences here.
The $E$ is trivially obtained using eq. 26 and $V^{r}_{ij}$. The only non-conservative force for the modified Active Ising model is the self-propulsion. Thus, $\langle\dot{\Sigma}^{nr}\rangle=\langle\dot{\Sigma}^{ch}\rangle=0$. The macroscopic mean EPR due to the self-propulsion reads:
@@
The macroscopic second law of thermodynamics reads $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+\langle\dot{\Sigma}^{sp}\rangle+d_{t}S^{gb}$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{sp}\rangle$. Other important implications for the phase diagram are reported in Ref. [86].
@@ B Non-reciprocal two species
We aim to study a more sophisticated model that combines non-reciprocity and self-propulsion simultaneously. The bird-flocking phenomena for a predator-prey bird species with two different preferred flying directions are denoted by $+$ and --. The predator (prey) is a Falcon (Starling). The microscopic interaction parameters for this model are $v_{s^{+}s^{+}}=v_{s^{-}s^{-}}=v_{f^{+}f^{+}}=v_{f^{-}f^{-}}=-1$ and, $v_{s^{+}s^{-}}=v_{f^{+}f^{-}}=1$ and, $v_{s^{+}f^{+}}=v_{s^{-}f^{-}}=v_{s^{+}f^{-}}=v_{s^{-}f^{+}}=1$ and, $v_{f^{+}s^{+}}=v_{f^{-}s^{-}}=v_{f^{+}s^{-}}=v_{f^{-}s^{+}}=-1$. It leads to the macroscopic interaction coefficients, $V_{s^{+}s^{+}}^{r}=V_{s^{-}s^{-}}^{r}=V_{f^{+}f^{+}}^{r}=V_{f^{-}f^{-}}^{r}=-\beta+\frac{\beta^{2}}{2\Omega}$, $V_{s^{+}s^{-}}^{r}=V_{f^{+}s^{-}}^{r}=\beta+\frac{\beta^{2}}{2\Omega}$, $V_{s^{+}f^{+}}^{r}=V_{s^{+}f^{-}}^{r}=V_{s^{-}f^{+}}^{r}=V_{s^{-}f^{-}}^{r}=\frac{\beta^{2}}{2\Omega}$, $V_{s^{+}f^{+}}^{nr}=V_{s^{+}f^{-}}^{nr}=V_{s^{-}f^{+}}^{nr}=V_{s^{-}f^{-}}^{nr}=\beta$. $E$ is trivially obtained using eq. 26 and $V^{r}_{ij}$. The macroscopic self-propulsion EPR reads:
$\begin{split}\langle\dot{\Sigma}^{sp}\rangle&=\int_{\mathcal{V}}2\bigg{[}F_{s}^{sp}\sinh{\left(\frac{F_{s}^{sp}}{2}\right)}\left(D_{s^{+}}+D_{s^{-}}\right)\\ &\hskip 42.67912pt+F_{f}^{sp}\sinh{\left(\frac{F_{f}^{sp}}{2}\right)}\left(D_{f^{+}}+D_{f^{-}}\right)\bigg{]}.\end{split}$ (69)
The macroscopic non-reciprocal EPR reads:
Here, $\rho_{s}=\rho_{s^{-}}+\rho_{s^{+}}$ and $\rho_{f}=\rho_{f^{-}}+\rho_{f^{+}}$. The set of eqs. 69 and 70 gives the macroscopic second law of thermodynamics, $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+\langle\dot{\Sigma}^{nr}\rangle+\langle\dot{\Sigma}^{sp}\rangle+d_{t}S^{gb}$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle+\langle\dot{\Sigma}^{sp}\rangle$, in the absence of the external driving of $E$.
6 Conclusion and Outlook
We have formulated a generic novel framework of stochastic thermodynamics for non-reciprocal systems that relies on systematic thermodynamically consistent coarse-graining approaches. Hence, the applicability of stochastic thermodynamics is broadened across different observation scales of the system. We further decompose the mean EPR into four orthogonal contributions, namely conservative, non-reciprocal, external chemical/self-propulsion driving and the rate of change in free energy (driving work). They correspond to entropy production cost associated with the relaxation (towards the reference Boltzmann distribution), sustaining the vorticity currents, sustaining the dissipative transition currents and the quasistatic mean stochastic work, respectively. Importantly, the systematic coarse-graining using the Doi-Peliti field theory ensures the equivalence of the systems’ dynamics and thermodynamics across the observation scale. LDB provides a thermodynamically consistent formulation across the scales.
We compute the dynamic equations of motion for the macrostate using the Langevin approximation, which exhibits multiplicative demographic noise. This generalizes the Macroscopic Fluctuation Theory. We demonstrated that the microscopic non-reciprocal interactions lead to the manifestation of Onsager’s non-reciprocal relations on the macroscale. We formulate the fluctuation response relation and its generalizations, namely higher-order current cumulant response relations. In the spirit of stochastic thermodynamics, we formulate the generic fluctuation relations for non-reciprocal systems. We obtain the tightest Thermodynamic Kinetic Uncertainty Relation for the non-reciprocal systems. It relies on using the observable transition currents and the observable vorticity that lies in the transition and state space respectively. Our framework opens up a plethora of natural extensions. In appendix B, we briefly highlight how to incorporate the information thermodynamics for non-reciprocal systems. Different aspects of this work are summarized in table 3.
@@ 6 Conclusion and Outlook
Microscopic section 2 Multiparticle state probability $P_{\{N\}}$ Master equation eq. 7 eq. 5 $s_{state}$ in eq. 9 $s_{b}$ in eq. 1 $\epsilon_{i}^{r}$ in eq. 1 $f_{i}^{nr}$ in eq. 1 $f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{f}_{i}^{sp}$ in eq. 5
Macroscopic section 3 Fluctuating macrostate density $\rho_{i}$ Generalized macroscopic fluctuation theory eq. 28 eq. 27 $S_{state}$ in eq. 34 $\mu_{i}^{id}=\ln{(\rho_{i})}$ $\mu_{i}^{r}$ in eq. 24 $F_{i}^{r}$ in eq. 24 $F_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{F}_{i}^{sp}$ in eq. 27
Hydrodynamic order parameters [64, 65] Phenomenological order parameters $\phi=g(\rho_{i})$ Model A, B, etc. Absent $\mu_{\phi_{i}}^{id}=\frac{\delta\mathcal{F}_{GL}^{id}}{\delta\phi_{i}}$ $\mu_{\phi_{i}}=\frac{\delta\mathcal{F}_{GL}}{\delta\phi_{i}}$ Usually absent but considered in [57, 58, 59, 60, 61, 62, 63, 76, 77, 78] Usually absent but considered in [87, 139, 140]
@@
Multi-body microscopic interactions. —  One can explore the consequences of the nonlinearity dependence of $\epsilon_{i}$ on $N_{i}$, which exhibits a richer phenomenology [59]. The structure of our framework is robust to such a modification to the non-linear dependence of Boltzmann weights on the particle number. Hence, a straightforward generalization is obtained [86]. Importantly, the higher-order multi-body interaction plays a key role in ensuring that the interaction energy functional is bounded from below.
@@
Linear cyclic CRN. —  A recent development in the interacting (non-ideal) CRN has led to the applicability of the methodology developed for non-interacting CRN [1, 264, 223, 265, 225, 266, 267] to the interacting CRN under certain conditions [268]. We intentionally avoided the interplay of the interacting systems and the topological properties of the CRN encapsulating the underlying reactive transitions. Nevertheless, the interacting CRNs are fundamentally similar to the noninteracting CRNs, provided that certain constraints on the topological properties of the CRN are met. An important implication is vanishing diffusive currents [268]. In particular, the interaction of diffusive and reactive currents has been the key mechanism of pattern formation [269]. This opens up an interesting avenue to explore, under a more generic interacting chemical reaction network that does not satisfy the constraints in Ref. [268]. In particular, the study of the interplay of the diffusive fluxes and the chemical reaction network and its thermodynamic and dynamic (pattern formation) implications. We hypothesize that the non-reciprocal systems might play a crucial role.
@@
Non-Linear CRN. —  Our framework can be straightforwardly incorporated into the non-linear deterministic CRN [151, 270] by incorporating the topological structure of the CRN. Moreover, the stochastic non-linear CRN exhibits more sophisticated effects arising due to the interplay of the non-linearity and stochasticity [271, 272, 273, 274, 275, 276]. It is an interesting avenue to explore the non-linear stochastic CRN by incorporating non-reciprocal interactions. This is because our framework systematically incorporates the microscopic occupancy fluctuations on the macroscale.
@@
The role of the demographic noise. —  Demographic noise plays a key role in understanding the physics of ecological models [277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 287, 288, 289]. For instance, steady state selection, stability of the attractors for the deterministic and stochastic dynamics [278, 279, 280, 281, 282, 283, 284, 285, 286, 287, 288, 289] and the mismatch between the deterministic and stochastic dynamics [290, 291, 292, 136]. It is an interesting avenue to explore the model-specific implications of demographic noise and potentially richer novel phenomena using eq. 28. Importantly, our framework enables us to compute the thermodynamic dissipation cost associated with different phases.

### v3

@@ B Predator-Prey Model
We consider a predator-prey system with a minimalist attraction-repulsion mechanism, such as dogs-sheep, without any self-propulsion. Thus, the microscopic interaction rules are $v_{ds}=-1$, $v_{sd}=1$, $v_{dd}=0$, $v_{ss}=-1$, which implies $v_{ds}^{r}=-1$, $v_{sd}^{r}=1$, $v_{dd}^{r}=0$, $v_{ss}^{r}=-1$. This leads to the macroscopic interaction rules $V_{sd}^{r}=\frac{\beta^{2}}{2\Omega}$, $V_{sd}^{nr}=\beta$, $V_{ds}^{nr}=-\beta$, and $V_{ss}^{r}=-\beta+\frac{\beta^{2}}{2\Omega}$. We have intentionally chosen $v_{ss}=-1$, which models an attractive interaction between the sheep, physically corresponding to the flocking of the sheep herd. In addition, $f_{ds}^{ch}=\vec{f}_{s}^{sp}=\vec{f}_{d}^{sp}=0$. Then, the macroscopic coarse-grained energy functional for this model obtained using the density of the sheep ($\rho_{s}$) and the dogs ($\rho_{d}$) reads:
@@
where the reciprocal macroscopic Boltzmann weights are $\mu_{d}^{r}=\ln{\rho_{d}}+V_{sd}^{r}\rho_{s}$ and $\mu_{s}^{r}=\ln{\rho_{s}}+V_{sd}^{r}\rho_{d}+V_{ss}^{r}\rho_{s}$, and the non-reciprocal macroscopic Boltzmann weights are $F_{d}^{nr}=V_{ds}^{nr}\rho_{s}$, $F_{s}^{nr}=V_{sd}^{nr}\rho_{d}$. The diffusive mobilities are then $D_{d}=\tilde{d}_{d}\rho_{d}e^{(V_{ds}^{r}+V_{ds}^{nr})\rho_{s}}$ and $D_{s}=\tilde{d}_{s}\rho_{s}e^{\left(V_{ss}^{r}\rho_{s}+(V_{sd}^{r}+V_{sd}^{nr})\rho_{d}\right)}$. In addition, $\langle\dot{\Sigma}^{sp}\rangle=\langle\dot{\Sigma}_{\mathcal{R}}^{nr}\rangle=\langle\dot{\Sigma}^{ch}\rangle=0$, since $F_{ds}^{ch}=\vec{F}_{s}^{sp}=\vec{F}_{d}^{sp}=0$. The expression for $\langle\dot{\Sigma}^{nr}\rangle$ reads:
@@
Assuming $\langle W\rangle=0$, the second law of thermodynamics for this model is $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+d_{t}S^{gb}+\langle\dot{\Sigma}^{nr}\rangle$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle$.
The change in the symmetric part of the macrostate correlations is $O(1/\Omega)$, as $V_{sd}^{nr}\propto O(1/\Omega)$, compared to the chemo-sensing bacteria in section A, where $V_{sd}^{nr}\propto O(1)$, which highlights a very important difference between the two models. Importantly, despite the similarity of $\langle\dot{\Sigma}^{nr}\rangle$ for both models, the symmetric part of the macrostate correlations distinguishes between the underlying different physical mechanisms that could generate similar vorticity currents. In the chemo-sensing bacteria section, it is one-way attraction, and here it is a mutual attraction-repulsion mechanism between the dogs and sheep that leads to non-reciprocal vorticity currents.
@@ A Single species
Here, we consider the thermodynamically consistent Active Ising Model (AIM). The AIM consists of two different types of self-propelled particles, positive and negative [68]. The particles of the same type attract each other, otherwise repel. Hence, the microscopic interaction rules are $v_{++}=-1$, $v_{--}=-1$, and $v_{-+}=1$. This leads to the macroscopic interaction coefficients $V_{--}^{r}=V_{++}^{r}=-1+\frac{\beta}{2\Omega}$ and $V_{-+}^{r}=1+\frac{\beta}{2\Omega}$. The macroscopic reciprocal Boltzmann weights are $\mu_{+}^{r}=\log{\rho_{+}}+\beta\left(V_{++}^{r}\rho_{+}+V_{+-}^{r}\rho_{-}\right)$ and $\mu_{-}^{r}=\log{\rho_{-}}+\beta\left(V_{--}^{r}\rho_{-}+V_{-+}^{r}\rho_{+}\right)$. The diffusive mobility for the density fields $\rho_{+}$ and $\rho_{-}$ reads $D_{+}=\tilde{d}e^{\mu_{+}^{r}}$ and $D_{-}=\tilde{d}e^{\mu_{-}^{r}}$, respectively. We report a detailed study in Ref. [72, 87] and briefly outline the thermodynamic consequences here.
The macroscopic energy functional $E$ is trivially obtained using eq. 26 and $V^{r}_{ij}$. The only non-conservative force for the thermodynamically consistent AIM is the self-propulsion. Thus, $\langle\dot{\Sigma}^{nr}\rangle=\langle\dot{\Sigma}^{ch}\rangle=0$. The macroscopic mean EPR due to the self-propulsion reads:
@@
The macroscopic second law of thermodynamics reads $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+\langle\dot{\Sigma}^{sp}\rangle+d_{t}S^{gb}$, or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{sp}\rangle$. We report other important implications for the correctness of the macroscopic phase diagram in Ref.[87].
@@ B Non-reciprocal two species
Microscopic description Multiparticle state probability $P_{\{N\}}$ Master equation eq. 7 eq. 5 $s_{state}$ in eq. 9 $s_{b}$ in eq. 6 $\epsilon_{i}^{r}$ in eq. 1 $f_{i}^{nr}$ in eq. 1 $f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{f}_{i}^{sp}$ in eq. 5
Macroscopic fluctuating description Fluctuating macrostate density $\rho_{i}$ Generalized macroscopic fluctuation theory eq. 29 eq. 28 $S_{state}$ in eq. 35 $\mu_{i}^{id}=\ln{(\rho_{i})}$ $\mu_{i}^{r}$ in eq. 24 $F_{i}^{nr}$ in eq. 24 $F_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{F}_{i}^{sp}$ in eq. 28
Hydrodynamic order parameters [64, 65] Phenomenological order parameters $\phi=g(\rho_{i})$ Model A, B, etc. Absent $\mu_{\phi_{i}}^{id}=\frac{\delta\mathcal{F}_{GL}^{id}}{\delta\phi_{i}}$ $\mu_{\phi_{i}}=\frac{\delta\mathcal{F}_{GL}}{\delta\phi_{i}}$ Usually absent but considered in [57, 58, 59, 60, 61, 62, 63, 76, 77, 78, 79] Usually absent but considered in [88, 145, 146]
@@
We aim to study a more sophisticated model that combines both non-reciprocity and self-propulsion simultaneously. The bird-flocking phenomena for a predator-prey bird species with two different preferred flying directions are denoted by ++ and -. The predator (prey) is a Falcon (Starling). The microscopic interaction parameters for this model are $v_{s^{+}s^{+}}=v_{s^{-}s^{-}}=v_{f^{+}f^{+}}=v_{f^{-}f^{-}}=-1$ and $v_{s^{+}s^{-}}=v_{f^{+}f^{-}}=1$ and $v_{s^{+}f^{+}}=v_{s^{-}f^{-}}=v_{s^{+}f^{-}}=v_{s^{-}f^{+}}=1$ and $v_{f^{+}s^{+}}=v_{f^{-}s^{-}}=v_{f^{+}s^{-}}=v_{f^{-}s^{+}}=-1$. It leads to the macroscopic interaction coefficients $V_{s^{+}s^{+}}^{r}=V_{s^{-}s^{-}}^{r}=V_{f^{+}f^{+}}^{r}=V_{f^{-}f^{-}}^{r}=-\beta+\frac{\beta^{2}}{2\Omega}$, $V_{s^{+}s^{-}}^{r}=V_{f^{+}s^{-}}^{r}=\beta+\frac{\beta^{2}}{2\Omega}$, $V_{s^{+}f^{+}}^{r}=V_{s^{+}f^{-}}^{r}=V_{s^{-}f^{+}}^{r}=V_{s^{-}f^{-}}^{r}=\frac{\beta^{2}}{2\Omega}$, $V_{s^{+}f^{+}}^{nr}=V_{s^{+}f^{-}}^{nr}=V_{s^{-}f^{+}}^{nr}=V_{s^{-}f^{-}}^{nr}=\beta$. The macroscopic interaction energy functional $E$ is trivially obtained using eq. 26 and $V^{r}_{ij}$. The macroscopic self-propulsion EPR reads:
$\begin{split}\langle\dot{\Sigma}^{sp}\rangle&=\int_{\mathcal{V}}2\bigg[F_{s}^{sp}\sinh{\left(\frac{F_{s}^{sp}}{2}\right)}\left(D_{s^{+}}+D_{s^{-}}\right)\\ &\hskip 42.67912pt+F_{f}^{sp}\sinh{\left(\frac{F_{f}^{sp}}{2}\right)}\left(D_{f^{+}}+D_{f^{-}}\right)\bigg],\end{split}$ (72)
and the macroscopic non-reciprocal EPR reads:
Here, $\rho_{s}=\rho_{s^{-}}+\rho_{s^{+}}$ and $\rho_{f}=\rho_{f^{-}}+\rho_{f^{+}}$ quantify the total density of the Starling and Falcon, irrespective of their flying direction. The set of eqs. 72 and 73 gives the orthogonal decomposition of the macroscopic second law of thermodynamics, $\langle\dot{\Sigma}\rangle=-d_{t}\langle E\rangle+\langle\dot{\Sigma}^{nr}\rangle+\langle\dot{\Sigma}^{sp}\rangle+d_{t}S^{gb}$. Or equivalently, $\langle\dot{\Sigma}\rangle=-d_{t}D^{KL}_{E}(\mathcal{P}[\{\rho\}])+\langle\dot{\Sigma}^{nr}\rangle+\langle\dot{\Sigma}^{sp}\rangle$, in the absence of the external driving of $E$, the driving work is $\langle W\rangle=0$.
6 Conclusion and Outlook
We have formulated a generic novel framework of stochastic thermodynamics for non-reciprocal systems that relies on systematic thermodynamically consistent coarse-graining approaches. The novel formulation of the Local Detailed Balance condition for non-reciprocal systems provides a thermodynamically consistent formulation across the scales. Hence, our framework broadens the applicability of stochastic thermodynamics to different observation scales of the system. We further decompose the mean EPR into four orthogonal contributions, namely the conservative, the non-reciprocal, the external chemical/self-propulsion driving, and the rate of change in the free energy (driving work). They correspond to the entropy production cost associated with the relaxation (towards the reference Boltzmann distribution), sustaining the vorticity currents (due to non-reciprocity), sustaining the dissipative transition currents, and the quasistatic mean stochastic work, respectively. Importantly, the systematic coarse-graining using the Doi-Peliti field theory ensures the equivalence of the systems’ dynamics and thermodynamics across the observation scale.
We compute the dynamic equations of motion for the macrostate using the Langevin approximation, which exhibits the multiplicative demographic noise. This generalizes the Macroscopic Fluctuation Theory to non-reciprocal and externally driven systems. We demonstrated that the microscopic non-reciprocal interactions lead to the manifestation of Onsager’s non-reciprocal relations on the macroscale. We formulate the fluctuation-response relation and its generalizations, namely the higher-order current cumulant response relations. In the spirit of stochastic thermodynamics, we formulate the generic fluctuation relations for non-reciprocal systems. We obtain the tightest Thermodynamic Kinetic Uncertainty Relation for the non-reciprocal systems. It relies on using the observable transition currents and the observable vorticity that lies in the transition and state space, respectively. We briefly highlight how to incorporate the information thermodynamics for non-reciprocal systems. Our framework opens up a plethora of model-specific extensions of non-reciprocal systems across a different observable scale. Different aspects of this work are summarized in table 5.
@@ 6 Conclusion and Outlook
Multi-body microscopic interactions. —  One can explore the consequences of the nonlinear dependence of $\epsilon_{i}$ on $N_{i}$, which exhibits a richer phenomenology [59]. The structure of our framework (both dynamic and thermodynamic) is robust to such a modification to the non-linear dependence of the Boltzmann weights on the particle number. Hence, a straightforward generalization is obtained [87]. Importantly, the higher-order multi-body interaction plays a key role in ensuring that the interaction energy functional is bounded from below.
@@
Linear cyclic CRN. —  A recent development in the interacting (non-ideal) CRN has led to the applicability of the methodology developed for the non-interacting CRN [1, 274, 229, 275, 231, 276, 277] to the interacting CRN under certain conditions [278]. We intentionally avoided the interplay of the interacting systems and the topological properties of the CRN encapsulating the underlying reactive transitions. Nevertheless, the interacting CRNs are fundamentally similar to the non-interacting CRNs, provided that certain constraints on the topological properties of the CRN are met. An important implication is the vanishing of diffusive currents [278]. However, the interaction of diffusive and reactive currents has been the key mechanism of pattern formation [279]. Therefore, this opens up an interesting avenue to explore, under a more generic condition of a spatially extended interacting CRN that does not satisfy the constraints in Ref.[278]. The study of the interplay between diffusive fluxes and the reactive fluxes modeled by interacting CRNs and their thermodynamic and dynamic implications for pattern formation remains an open problem. We expect that the spatially extended non-reciprocal systems might exhibit a richer phenomenology.
@@
Nonlinear CRN. —  Our framework can be straightforwardly incorporated into the non-linear deterministic CRN [157, 280] by incorporating the topological structure of the CRN. Moreover, the stochastic non-linear CRN exhibits more sophisticated effects arising due to the interplay of non-linearity and stochasticity [281, 282, 283, 284, 285, 286]. It is an interesting avenue to explore the non-linear stochastic CRN by incorporating non-reciprocal interactions. Since our framework systematically incorporates the microscopic occupancy and transition fluctuations on the macroscale, it is an inherently better starting point than the deterministic CRN to study such effects.
@@
The role of demographic noise. —  Demographic noise plays a key role in understanding the physics of ecological models [287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299]. For instance, the steady-state selection, the stability of the attractors for the deterministic and stochastic dynamics [288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299], and the mismatch between the deterministic and stochastic dynamics [300, 301, 302, 142]. It is an interesting avenue to explore the model-specific implications of demographic noise and potentially richer novel phenomena using the exact eq. 29 for their thermodynamically consistent dynamics. Importantly, our framework enables us to compute the thermodynamic dissipation cost associated with the different phases as well.
The finite-time optimal control of non-reciprocal systems —  We have omitted the time-dependent change of the control parameters of the non-reciprocal models, except for the slow driving of the control parameters of the conservative energy functional $E$ that quantifies the quasistatic work. However, a novel framework for the finite-time optimal control of any control parameters of far-from-equilibrium systems has been recently developed in Ref.[303]. It has revealed the importance of the exact non-quadratic dissipation functions and the discontinuous endpoint jumps in optimal driving protocols, which have been shown to be of paramount importance for far-from-equilibrium systems [89, 90]. It is an interesting avenue to explore model-specific phenomena for the optimal control of non-reciprocal systems in a finite driving time.